OpenStudy (anonymous):

Tell whether each function is linear. If so, graph the function . 2x + 3y = 5 can someone show me how to do it please

OpenStudy (abb0t):

Firstly, this is Chemistry section, not math. If this IS for a chemistry course, I think you should brush up on your basic algebra as it would be quite useful when doing experiments and making graphs. With all that being said, You should recognize the function of a line: y = mx+b use your algebra tools to re-write the function in that form where m = slope b = y-coordinate.

OpenStudy (kittykitty04):

Step A) 2x + 3y = 5 Step B) Move the 2x over to the right side by subtracting it so that the equation can look like y=mx+b - It should now look like 3y= -2x+5 Step C) Now you need to get rid of the 3 on the 'y' and that is by diving it by everything. The equation should now look like this :  <- not sure if you can see the picture there. I'll just type it out just incase. y= -2/3 +5/3. This is a linear line on the graph with the y coordinate starting at (0, 5/3) or known as (0,1.666) and it's slope is negative : -2/3. Hope this helps.

OpenStudy (anonymous):

thank u so much